How much does a sleeping beg cost? Lets say you want a sleeping bag that should keep you warm in temperatures frem range is given below. Assume that the population of x values has an approximately normal distribution. i 20°F to 45F. A random sample of prices (8) for sieeping bags in ehes temperature 45 55 90 120 0 90 30 23 100110 10595 105 60110 12095 90 60 () Use a cakculator with -$ 81.15 -$ 29.72 mean and sample standard deviation keys to find the saeple mean price and smple standard deviation s . ound your answers te two decmal pliaces) (b) Using the given data as representative of the answers to two decimal places ower imit 68 40 upper lima 90.80 population of prices or all summer sleeping bags, nnd a 90% medence interven he mean pre μ ofat-seer- y bag. (Round you Need Help? oa
Answer #1

Ans:

b)df=20-1=19

critical t value=tinv(0.1,19)=1.729

90% confidence interval for population mean

=81.15+/-1.729*(29.72/sqrt(20))

=81.15+/-11.49

=(69.66, 92.64)